# A Moreau Envelope Approach for LQR Meta-Policy Estimation

### v1

@@ IV-A Uncertain linear system
We used the following model parameters for the experiments in (1) and (2). The uncertainties in the system are $a\in[-1,1]^{2}$ and $b\in[-1,1]^{2}$. The matrices generating the system are $\mathcal{A}_{0}=\begin{bmatrix}0.7&-0.3&0.0&0.1\\ 0.5&-0.4&0.3&0.0\\ 0.0&0.4&0.2&-0.1\\ 0.2&0.0&0.4&0.6\end{bmatrix}$, $\mathcal{A}_{1}$ is a lower triangular matrix with all entries as 0.1, and $\mathcal{A}_{2}$ is a similar upper triangular matrix,
$\mathcal{B}_{0}=\begin{bmatrix}0.3&0.2\\ 0.1&0.5\\ 0.4&0.1\\ 0.0&0.1\end{bmatrix},\mathcal{B}_{1}=\begin{bmatrix}0.1&0.0\\ 0.0&0.1\\ 0.1&0.0\\ 0.0&0.1\end{bmatrix},\mathcal{B}_{2}=\begin{bmatrix}0.0&0.1\\ 0.1&0.0\\ 0.0&0.1\\ 0.1&0.0\end{bmatrix}$
@@
Four system realizations were generated using the parameters provided above. The MEMLQR algorithm was implemented with $S=300$, $P=2$, $\alpha=0.1$, $\beta=1$ and $\lambda=0.2$. To solve the optimization problem in Step 5, we use the model-based gradient descent algorithm from [3]. We show the convergence characteristics of the algorithm by plotting the Moreau regularized cost $(C_{\lambda}\bigl{(}K^{s}\bigr{)})$ against the number of outer iterations $(s)$ in Fig. 1(a). Further, in Fig. 1(b) we plot the accuracy $\mathopen{}\mathclose{{}\left(1-{|C_{z}(K^{N})-C_{z}(K^{n})|}/{C_{z}(K^{N})}}\right)$ for the cost incurred while adapting to a new system realization $z$, by the initializing a policy that minimizes the cumulative cost and the MEMLQR policy generated previously. It can be observed that the MEMLQR policy adapts faster to the optimal policy of the unseen realization $z$. We also plot in Fig. 1(c) the state and input trajectories of the unseen system realization $z$ for the policy generated when a policy gradient algorithm is initialized at the MEMLQR policy.
Figure 1: Depiction of properties of Algorithm 1: (a) Convergence of the MEMLQR algorithm: evolution of the Moreau envelope regularized cost $C_{\lambda}(K^{s})$ as the number of outer iterations $(s)$ increase. (b) The evolution of accuracy $\mathopen{}\mathclose{{}\left(1-{|C_{z}(K^{N})-C_{z}(K^{n})|}/{C_{z}(K^{N})}}\right)$ of the policy generated for an unseen realization $z$ by a model-based policy gradient framework when initialized using a policy obtained by minimizing the total LQR cost $C(\cdot)$ and by using the $K^{S}$ generated by the MEMLQR algorithm. Note that, the cost calculation here is over 50 randomly initial states. (c) State and input trajectories of the unseen realization $z$ for the estimate of the optimal policy $(K^{N})$ generated after $N=250$ iterations of the model-based policy gradient algorithm, which was initialized at $K^{S}$ generated by the MEMLQR algorithm.
@@ IV-B Comparison to the MAML approach
To empirically compare the performance of the MAML-based approach to that of ours, we borrowed the example presented in [20]. We generated ten different system realizations and obtained meta policies using the MAML-LQR algorithm from [20] and our MEMLQR algorithm. These policies are evaluated by comparing the convergence characteristics while performing model-free policy gradient for three previously unseen system realizations. It can be observed from Fig. 2 that at initialization, the MEMLQR policy has a lower cost than the MAML-LQR policy for all three realizations, thus leading to a faster improvement in the cost incurred.
Figure 2: Convergence to the optimal policy after initialization using the policy generated by MAML-LQR and MEMLQR for three random system realizations in a model-free setting. It can be observed that the cost incurred by policy generated by the MEMLQR approach is closer to the optimal value initially, thus aiding in faster convergence.
@@ V Conclusion
In this article, we explored personalization using Moreau envelopes to obtain an initialization for policy gradient algorithms (model-based/free) when dealing with uncertain linear systems, given that there is access to a finite number of uncertainty realizations. An iterative first-order framework, Algorithm 1, was presented to optimize this personalized cost function in a client-server setup, and its analysis showed the algorithm’s convergence to a first-order stationary point. Numerical experiments were provided to empirically show the advantage of designing and using the proposed framework’s meta-policy rather than the total cost for the same purpose. We also compared the results of our approach to that of an approach based on the MAML framework and empirically showed that our approach incurred a lower cost for unseen scenarios. As an avenue for future work, we plan to explore personalization methods to compute optimal policies for uncertain linear systems when the system parameters are distributed over nodes with limited communication capabilities.

### v2

@@ IV-A Uncertain linear system
We used the following model parameters for the experiments in (1). The uncertainties in the system are $a\sim\text{Unif}(-1,1)^{2}$ and $b\sim\text{Unif}(-1,1)^{2}$, where $\text{Unif}(-1,1)$ represents a uniform distribution between $-1$ and $1$. The matrices generating the system are $\small\mathcal{A}_{0}=\begin{bmatrix}0.7&-0.3&0.0&0.1\\ 0.5&-0.4&0.3&0.0\\ 0.0&0.4&0.2&-0.1\\ 0.2&0.0&0.4&0.6\end{bmatrix}$, $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ are lower and upper triangular matrices, respectively, with all entries as 0.1 and
$\small\mathcal{B}_{0}=\begin{bmatrix}0.3&0.2\\ 0.1&0.5\\ 0.4&0.1\\ 0.0&0.1\end{bmatrix},\mathcal{B}_{1}=\begin{bmatrix}0.1&0.0\\ 0.0&0.1\\ 0.1&0.0\\ 0.0&0.1\end{bmatrix},\mathcal{B}_{2}=\begin{bmatrix}0.0&0.1\\ 0.1&0.0\\ 0.0&0.1\\ 0.1&0.0\end{bmatrix}.$
@@
Four system realizations were generated using the parameters provided above. The MEMLQR algorithm was implemented with $S=300$, $P=2$, $\alpha=0.1$, $\beta=1$ and $\lambda=0.2$. To solve the optimization problem in Step 5, we use the model-based gradient descent algorithm from [3]. We show the convergence characteristics of the algorithm by plotting the Moreau regularized cost $(C_{\lambda}\bigl{(}K^{s}\bigr{)})$ against the number of outer iterations $(s)$ in Fig. 1(a). Note that we slightly abuse our notations here by representing the cost incurred while adapting to a new system realization $z$ as $C_{z}$. In Fig. 1(b) we plot the accuracy $\mathopen{}\mathclose{{}\left(1-{|C_{z}(K^{N})-C_{z}(K^{n})|}/{C_{z}(K^{N})}}\right)$ for the cost incurred while adapting to a new system, by initializing using the policy that minimizes the cumulative cost and the MEMLQR policy generated previously. It can be observed that the MEMLQR policy adapts faster to the optimal policy of the unseen realization $z$. We also plot in Fig. 1(c) the state and input trajectories of the unseen system realization $z$ for the policy generated when a policy gradient algorithm is initialized at the MEMLQR policy.
Figure 1: Depiction of properties of Algorithm 1: (a) Convergence of the MEMLQR algorithm: evolution of the Moreau envelope regularized cost $C_{\lambda}(K^{s})$ as the number of outer iterations $(s)$ increase. (b) The evolution of accuracy $\mathopen{}\mathclose{{}\left(1-{|C_{z}(K^{N})-C_{z}(K^{n})|}/{C_{z}(K^{N})}}\right)$ of the policy generated for an unseen realization $z$ by a model-based policy gradient framework when initialized using a policy obtained by minimizing the total LQR cost $C(\cdot)$ and by using the $K^{S}$ generated by the MEMLQR algorithm. Note that the cost calculation here is over 50 randomly initial states. (c) State and input trajectories of the unseen realization $z$ for the estimate of the optimal policy $(K^{N})$ generated after $N=250$ iterations of the model-based policy gradient algorithm, which was initialized at $K^{S}$ generated by the MEMLQR algorithm.
@@ IV-B Comparison to the MAML approach
To empirically compare the performance of the MAML-based approach to that of ours, we borrowed the example presented in [21]. We generated ten different system realizations and obtained meta policies using the MAML-LQR algorithm from [21] and our MEMLQR algorithm with three different values, i.e., $\lambda\in\mathopen{}\mathclose{{}\left\{0.02,0.2,2}\right\}$. These policies are evaluated by comparing the convergence characteristics while performing model-free policy gradient for three previously unseen system realizations. It can be observed from Fig. 2 that at initialization, the MEMLQR policy has a lower cost than the MAML-LQR policy for all three realizations, thus leading to a faster improvement in the cost incurred.
Figure 2: Convergence to the optimal policy after initialization using the policy generated by MAML-LQR and MEMLQR ($\lambda=0.02,0.2,2$) for three random system realizations in a model-free setting. It can be observed that the cost incurred by policy generated by the MEMLQR approach is closer to the optimal value initially, thus aiding in faster convergence. Also, it may be noted that for higher values of $\lambda$, the cost incurred is lower.
@@ V Conclusion
In this article, we explored personalization using Moreau envelopes to obtain an initialization for policy gradient algorithms (model-based/free) when dealing with uncertain linear systems, given that there is access to a finite number of uncertainty realizations. An iterative first-order framework, Algorithm 1, was presented to optimize this personalized cost function in a client-server setup, and its analysis showed the algorithm’s convergence to a first-order stationary point. Numerical experiments were provided to empirically show the advantage of designing and using the proposed framework’s meta-policy rather than the total cost. We compared the results of our approach to that of an approach based on the MAML framework and empirically showed that our approach incurred a lower cost for unseen scenarios.